# Percolation in semicontinuum geometries

### v1

@@ III.2.1 Squares and Rectangles in 2D
The $Y$-intercept of the straight line fit gives the true threshold $\phi_{c}(\infty)$. We calculate the horizontal and vertical thresholds separately. Typical plots are shown in Fig. 8. For rectangles, the horizontal and vertical effective thresholds for finite system sizes will be different as a result of the anisotropy of the shape. However, as we can see from Fig. 8, both converge to a single threshold value in the infinite system limit. This threshold isotropy is similar to what is seen for the percolation of extended objects in lattices and continuum systems jasna ; klatt .
The percolation threshold values obtained from the simulations for squares of different sizes are listed in Table. 1 along with the excluded volume result based on Eq. (4). To obtain the numerical values of the thresholds from Eq. (4), we have assumed that the value of $B_{c}$ is the same as that of the corresponding continuum system ($B_{c}\approx 4.3953711$ koza ). The theoretical and simulation results are shown in Fig. 9 in which a comparison with the corresponding model on lattice system from koza is also shown. We can see that the threshold values increase monotonically and there is a very good agreement between the theoretical predictions and simulation results. The excellent agreement seen even for relatively small values of $k$ validates the assumption that the critical connectivity of rectangles is the same (i.e. value of $B_{c}$) for continuum and semicontinuum models of overlapping squares. We see similar good agreement between simulation results and excluded volume predictions for other shapes like disks and rectangles indicating the validity of this assumption in general.
Figure 8: $\phi_{c}^{eff}$ vs width $\Delta$ for horizontal and vertical percolation for $2\times 1$ rectangles in lanes. The best straight-line fits with the largest three system sizes are shown. The $y$- intercept gives the true thresholds in the thermodynamic limit. Here for horizontal percolation, the $y$-intercept is $\phi_{c}^{k_{1},k_{2}}=0.5314(4)$ and for vertical percolation it is $\phi_{c}^{k_{1},k_{2}}=0.5309(1)$.
@@
1 0.5310(3) 0.5193
2 0.5893(2) 0.5848
3 0.6157(15) 0.6101
4 0.6228(18) 0.6234
5 0.6323(5) 0.6317
6 0.6370(10) 0.6373
7 0.6404(12) 0.6414
8 0.6459(22) 0.6445
9 0.6496(12) 0.6469
10 0.6498(9) 0.6488
11 0.6521(15) 0.6504
12 0.6499(8) 0.6517
13 0.6527(1) 0.6529
14 0.6538(19) 0.6538
15 0.6547(7) 0.6547
20 0.6582(20) 0.6577
Figure 9: Variation of percolation threshold $\phi_{c}^{k}$ with $k$ for overlapping squares in lanes compared with that for overlapping squares on a square lattice given in koza . Simulation results and excluded volume results are shown in both the cases. Corresponding continuum value is also shown for comparison ($\phi_{c}^{\infty}\approx 0.6667$ mertens ). Semicontinuum model of squares shows a monotonic increase of threshold with $k$ whereas lattice models show a non-monotonic behaviour (See Ref. koza ).
@@
1 0.5310(1) 0.5310(6) 0.5310(3)
2 0.5309(1) 0.5314(4) 0.5312(2)
3 1 0.5331(2) 0.5310(10) 0.5325(1) 0.5193
4 05310(4) 0.5305(1) 0.5308(2)
5 0.5327(15) 0.5332(22) 0.5330(13)
1 0.5892(1) 0.5883(9) 0.5888(5)
2 0.5891(3) 0.5895(2) 0.5893(2)
3 2 0.5902(3) 0.5899(2) 0.5901(2) 0.5848
4 0.5897(2) 0.5902(10) 0.5899(5)
5 0.5885(4) 0.5892(3) 0.5889(3)
1 0.6135(29) 0.6122(30) 0.6129(21)
2 0.6135(9) 0.6168(8) 0.6152(6)
3 3 0.6156(12) 0.6158(28) 0.6157(15) 0.6101
4 0.6144(12) 0.6148(16) 0.6146(10)
5 0.6142(25) 0.6132(30) 0.6137(20)
We also evaluate the correlation length exponent $\nu$ for the semicontinuum model of rectangles. We use the scaling relation stauffer ,
@@
The slope of the $\log-\log$ plot between $\Delta(L)$ and $L$ gives the exponent $1/\nu$. A sample plot for the semicontinuum problem of $2\times 1$ rectangles is shown in Fig. 10. The obtained values for all systems involving rectangles of different sizes are found to be in line with the standard value of $41/\nu=3/$ 4 for 2D site percolation stauffer ; christensen . As a confirmation, we verify that a good scaling collapse is obtained for the percolation probability $P(\phi)$ with the determined values of the threshold and the correlation length exponent. $P(\phi)$ is expected to follow the scaling relation stauffer ,
@@
Figure 10: $\log-\log$ plot of the width of the percolation curves $\Delta$ with system size $L$ along with the best straight line fit for $2\times 1$ rectangles. The straight-line fit is obtained with the largest three system sizes. Slope of the straight line gives $1/\nu=0.74\pm 0.01$ (See Eq. (18)).
III.2.2 Other 2D models: Disks and parallelograms in lanes
In this section, we consider two other semicontinuum systems in 2D. The first one is a 2D semicontinuum model of disks of single radius described in Sec.III.1. Disks of diameter $k$ are placed randomly in lanes until a spanning cluster appears (see Fig. 3 (b)). The threshold values are obtained in a similar manner described in Sec.III.2.1. The percolation thresholds obtained for the diameter values of $k=2,3,4,5,6,7$ are listed in Table 3. We can see a decreasing trend in the threshold with increasing $k$ as predicted by Eq. (16). Again, excellent agreement is seen between the predicted values and the simulation results.
@@ III.2.2 Other 2D models: Disks and parallelograms in lanes
Table 3: Percolation threshold $\phi_{c}^{k}$ for semicontinuum disk models of different diameter values $k$ determined from simulations and Eq.(16). Percolation threshold obtained from simulation is the average of horizontal and vertical thresholds. Threshold calculated from Eq.(16) uses $B_{c}\approx 4.512348$ (taken from corresponding continuum model) xun .
The second shape we consider is overlapping parallelograms. We can view this system as the semicontinuum limit of overlapping parallelograms on a triangular lattice (Fig. 11). Here, we first verify a prediction made in Ref. jasna that on a triangular lattice, parallelograms of width $k_{2}=1$ will have a constant percolation threshold independent of its length. The numerical value of the threshold is given by $\phi_{c}=1-\exp(-B_{c}/6)$ where $B_{c}$ is expected to be the same as that of the continuum percolation problem of rectangles. Table. 4 lists the threshold values for widths $k_{2}=1$ and $2$. We can see that for width $k_{2}=2$, there is a clear increasing trend in the threshold values, whereas for $k_{2}=1$, there is no discernible trend and the threshold values are nearly constant within the numerical accuracy of our simulations. Note that $k_{1}=k_{2}=1$ corresponds to the usual site percolation on the triangular lattice for which the threshold is known exactly to be $1/2$  sykes .
Now, for the corresponding semicontinuum problem, we can imagine taking the triangular lattice and making the $X$ direction continuous while maintaining a discrete geometry in the $Y$ direction. We can see that with periodic boundary conditions applied, the semicontinuum problem of parallelograms of size $k_{1}\times k_{2}$ (note that here $k_{2}$ refers to the number of lanes the object covers in the vertical direction and not the Euclidean length) is the same as that of rectangles of the same size. Thus we expect the same thresholds for the two systems which we can verify by comparing Tables 2 and 5. We also expect that parallelograms in a continuum geometry have the same threshold as squares or rectangles in a continuum as the shapes are related to each other by a shearing transformation torquato ; modenov , which we verify in Table. 6.
@@
1 0.5007(2)
2 0.5160(13)
3 1 0.5189(19) 0.5193
4 0.5162(4)
5 0.5233(20)
6 0.5163(38)
1 0.5148(4) 0.5193
2 0.5466(41) 0.5503
3 2 0.5602(14) 0.5614
4 0.5691(20) 0.5671
5 0.5770(30) 0.5706
6 0.5767(21) 0.5729
Table 4: Percolation threshold $\phi_{c}^{k_{1},k_{2}}$ for parallelograms of width $k_{2}=1$ and width $k_{2}=2$ on triangular lattice from simulations and the excluded volume results jasna . Percolation threshold obtained from simulation is the average of horizontal and vertical thresholds.
@@
5 1 0.6685(40)
4 1 0.6664(30)
5 2 0.6656(18)
4 2 0.6652(16)
III.2.3 Cuboids in 3D
As an example of a semicontinuum system in 3D, we consider overlapping cuboids in lanes having continuous $X$ and $Y$ coordinates while having discrete $Z$ coordinates (see Fig.3 (a)). The thresholds are determined as in the 2D case and are given in Table. 7 along with the theoretical predictions based on Eq. (8). We can see that, in agreement with predictions of Eq. (8), threshold $\phi_{c}^{k_{1},k_{2},k_{3}}$ is independent of $k_{1}$ and $k_{2}$, for a given value of $k_{3}$. The numerical estimates of the thresholds are obtained by assuming $B_{c}\approx 2.5978$  koza to be that of the corresponding continuum problem.
$k_{1}$ $k_{2}$ $k_{3}$ $\phi_{c}^{x}$ $\phi_{c}^{y}$ $\phi_{c}^{z}$ $\phi_{c}$ Average $\phi_{c}$ theory (See Eq.(8), $B_{c}\approx 2.5978$ koza )
@@ IV Conclusion
We analyze a number of percolation problems of overlapping objects in 2D and 3D. In 2D, we considered the percolation of overlapping rectangles, disks, and parallelograms, and in 3D, we considered cuboids. We adapted the excluded volume argument to a semicontinuum setting and studied the dependence of the percolation threshold on the linear measures characterizing the basic percolation units considered. We show that in 2D, for rectangles, the percolation threshold is independent of the length (linear measure along the continuum direction) of the rectangles and is a constant for a given width (linear measure along the discrete direction), even when we have a distribution of lengths. This contrasts with the aspect-ratio-independent thresholds seen in continuum models. Aligned overlapping squares in lanes show a monotonic increase in threshold with increasing size of the squares. This is in contrast to the non-monotonic behavior for the threshold observed for overlapping squares on lattices koza . We analyzed overlapping disks and parallelograms in 2D semicontinuum geometries. For disks, we find that the threshold decreases monotonically with their radius.
@@
Models in lattice and continuum geometries have been extensively studied in several fields within Science, Engineering, and Mathematics. In many of them, studying how the properties of a system change as we go from a lattice to continuum geometry is a question of great significance. The semicontinuum geometry introduced here describes a system that lies between the two extremes of lattice and continuum geometries. We studied the specific problem of the percolation of overlapping objects in such geometries, but it is obvious that many other models may be considered in such geometries and could be of interest in many fields opening up directions for further investigations. A particular example is that of models with hardcore exclusion that have been extensively studied in lattices and show properties such as jamming and order-disorder transitions slutskii ; kundu .
Finally, we note that numerical values of the thresholds obtained here are based on the assumption that the connectivity at criticality is the same for similar shapes in continuum and semicontinuum systems, which allowed us to use the known values of $B_{c}$ from continuum systems. Our theoretical predictions of the numerical values in the models discussed turns out to be very accurate, justifying the assumption. However, theoretical treatments dedicated to semicontinuum geometries may be required to treat other percolation problems in such settings.
@@
Acknowledgements.

### v2

@@ 3.2.1 Squares and Rectangles in 2D:
The $Y$-intercept of the straight line fit gives the true threshold $\phi_{c}(\infty)$. We calculate the horizontal and vertical thresholds separately. Typical plots are shown in Fig. 8 (a). For rectangles, the horizontal and vertical effective thresholds for finite system sizes will be different as a result of the anisotropy of the shape. However, as we can see from Fig. 8 (a), both converge to a single threshold value in the infinite system limit. This threshold isotropy is similar to what is seen for the percolation of extended objects in lattices and continuum systems [25, 36].
The percolation threshold values obtained from the simulations for squares of different sizes are listed in Table. 1 along with the excluded volume result based on Eq. (4). To obtain the numerical values of the thresholds from Eq. (4), we have assumed that the value of $B_{c}$ is the same as that of the corresponding continuum system ($B_{c}\approx 4.3953711$ [20]). The theoretical and simulation results are shown in Fig. 9 in which a comparison with the results of the corresponding model on lattice system from [20] is also shown. We can see that the threshold values increase monotonically for the semicontinuum model and there is a very good agreement between the theoretical predictions and simulation results. (Note that the non-monotonicity observed for overlapping squares on a lattice in [20] is only due to a single point. However, the comparison of the case of overlapping cubes on a lattice [20] with the corresponding semicontinuum model also shows the same contrasting behavior between lattice and semicontinuum models (See Sec. 3.2.3 below)). The excellent agreement seen even for relatively small values of $k$ validates the assumption that the critical connectivity of rectangles is the same (i.e. value of $B_{c}$) for continuum and semicontinuum models of overlapping squares. We see similar good agreement between simulation results and excluded volume predictions for other shapes like disks and rectangles indicating the validity of this assumption in general.
Figure 8: (a) $\phi_{c}^{eff}$ vs width $\Delta$ for horizontal and vertical percolation for $2\times 1$ rectangles in lanes. The best straight-line fits with the largest five system sizes are shown. The $y$- intercepts give the true thresholds in the thermodynamic limit. Here for horizontal percolation, the $y$-intercept is $\phi_{c}^{k_{1},k_{2}}=0.53112(1)$ and for vertical percolation it is $\phi_{c}^{k_{1},k_{2}}=0.53114(2)$. (b) $\log-\log$ plot of the width of the percolation curves $\Delta$ with system size $L$ along with the best straight line fit for $2\times 1$ rectangles. The straight-line fit is obtained with the largest five system sizes. Slope of the straight line gives $1/\nu=0.74\pm 0.01$ (See Eq. (18)).
@@
    1 0.53111(3) 0.51932
    2 0.58909(3) 0.58483
    3 0.61230(3) 0.61010
    4 0.62476(4) 0.62347
    5 0.63269(3) 0.63173
    6 0.63807(5) 0.63735
    7 0.64184(8) 0.64141
    8 0.64477(5) 0.64449
    9 0.64721(3) 0.64690
    10 0.64909(7) 0.64884
    11 0.65061(8) 0.65044
    12 0.65200(4) 0.65177
    13 0.65308(2) 0.65290
    14 0.65416(8) 0.65387
    15 0.65482(3) 0.65471
    20 0.65785(4) 0.65769
[…]
@@
    1 0.53122(5) 0.53099(4) 0.53111(3)
    2 0.53114(2) 0.53112(1) 0.53113(1)
    3 1 0.53111(2) 0.53114(3) 0.53113(2) 0.51932
    4 053110(3) 0.53112(2) 0.53111(2)
    5 0.53111(4) 0.53119(2) 0.53115(2)
    1 0.58912(3) 0.58904(2) 0.58908(2)
    2 0.58909(4) 0.58909(3) 0.58909(3)
    3 2 0.58909(2) 0.58906(2) 0.58908(1) 0.58483
    4 0.58909(7) 0.58911(1) 0.58910(4)
    5 0.58904(2) 0.58918(9) 0.58911(5)
    1 0.61216(6) 0.61224(4) 0.61220(4)
    2 0.61233(2) 0.61224(5) 0.61229(3)
    3 3 0.61231(2) 0.61229(6) 0.61230(3) 0.61010
    4 0.61229(3) 0.61227(2) 0.61228(2)
    5 0.61228(2) 0.61231(3) 0.61230(2)
We also evaluate the correlation length exponent $\nu$ for the semicontinuum model of rectangles. We use the scaling relation [13],
@@
The slope of the $\log-\log$ plot between $\Delta(L)$ and $L$ gives the exponent $1/\nu$. A sample plot for the semicontinuum problem of $2\times 1$ rectangles is shown in Fig. 8 (b). The obtained values for all systems involving rectangles of different sizes are found to be in line with the standard value of $41/\nu=3/$ 4 for 2D site percolation [13, 14]. As a confirmation, we verify that a good scaling collapse is obtained for the percolation probability $P(\phi)$ with the determined values of the threshold and the correlation length exponent. $P(\phi)$ is expected to follow the scaling relation [13],
@@
3.2.2 Other 2D models: Disks and parallelograms in lanes
In this section, we consider two other semicontinuum systems in 2D. The first one is a 2D semicontinuum model of disks of single radius described in Sec.3.1. Disks of diameter $k$ are placed randomly in lanes until a spanning cluster appears (see Fig. 3 (b)). The threshold values are obtained in a similar manner described in Sec.3.2.1. The percolation thresholds obtained for the diameter values of $k=2,3,4,5,6,7$ are listed in Table 3. We can see a decreasing trend in the threshold with increasing $k$ as predicted by Eq. (16). Again, excellent agreement is seen between the predicted values and the simulation results.
Table 3: Percolation threshold $\phi_{c}^{k}$ for semicontinuum disk models of different diameter values $k$ determined from simulations and Eq.(16). Percolation threshold obtained from simulation is the average of horizontal and vertical thresholds. Threshold calculated from Eq.(16) uses $B_{c}\approx 4.512348$ (taken from corresponding continuum model) [42].
@@ 3.2.2 Other 2D models: Disks and parallelograms in lanes
The second shape we consider is overlapping parallelograms. We can view this system as the semicontinuum limit of overlapping parallelograms on a triangular lattice (Fig. 11). Here, we first consider a prediction made in Ref. [25] that on a triangular lattice, parallelograms of width $k_{2}=1$ will have a constant percolation threshold independent of its length. The numerical value of the threshold is given by $\phi_{c}=1-\exp(-B_{c}/6)$ where $B_{c}$ is expected to be the same as that of the continuum percolation problem of rectangles. Table. 4 lists the threshold values for widths $k_{2}=1,2$ and $3$. Note that $k_{1}=k_{2}=1$ corresponds to the usual site percolation on the triangular lattice for which the threshold is known exactly to be $1/2$  [43].
From Table. 4, we can see that for widths $k_{2}=2$ and $k_{2}=3$, there is a clear increasing trend in threshold and the values are in reasonable agreement with the predictions. However, an increasing trend in the threshold values is seen for width $k_{2}=1$ as well which shows a deviation from the prediction of a constant value. Thus the excluded volume theory prediction of a $k_{1}$-independent threshold doesn’t quite hold up in this case although from Fig. 10 (a), we can see that the percentage deviation of the thresholds from their asymptotic value is the lowest for width $k_{2}=1$ for any $k_{1}$. A similar comparison is shown for rectangles on square lattices in Fig. 10 (b). We can see that in this case, the percentage deviation of the threshold values from their asymptotic value is the smallest for width $k_{2}=2$ for any $k_{1}$. For the case of $k_{2}=2$, except for $k_{1}=1$, there is very little deviation (less than 1%) seen in the threshold values.
Figure 10: (a) Percentage deviation of the thresholds obtained from simulation from their approximate asymptotic values ($\phi_{c}^{\infty}$) for different $k_{1}$ for parallelograms of widths $k_{2}=1$, $k_{2}=2$ and $k_{2}=3$ on a triangular lattice. (b) Percentage deviation of the thresholds obtained from simulation from their approximate asymptotic values ($\phi_{c}^{\infty}$) for different $k_{1}$ for rectangles of widths $k_{2}=1$, $k_{2}=2$ and $k_{2}=3$ on a square lattice. The simulation data is an improvement over the one given in Ref. [25]. The approximate asymptotic values ($\phi_{c}^{\infty}$) are taken from Column 5 of Table. 2.
Now, for the corresponding semicontinuum problem, we can imagine taking the triangular lattice and making the $X$ direction continuous while maintaining a discrete geometry in the $Y$ direction. We can see that with periodic boundary conditions applied, the semicontinuum problem of parallelograms of size $k_{1}\times k_{2}$ (note that here $k_{2}$ refers to the number of lanes the object covers in the vertical direction and not the Euclidean length) is the same as that of rectangles of the same size. Thus we expect the same thresholds for the two systems which we can verify by comparing Tables 2 and 5. We also expect that parallelograms in a continuum geometry have the same threshold as squares or rectangles in a continuum as the shapes are related to each other by a shearing transformation [44, 45], which we verify in Table. 6.
@@
Table 4: Percolation threshold $\phi_{c}^{k_{1},k_{2}}$ for parallelograms of width $k_{2}=1$, $k_{2}=2$ and width $k_{2}=3$ on triangular lattice from simulations and the excluded volume results [25]. Percolation threshold obtained from simulation is the average of horizontal and vertical thresholds.
@@
    1 0.49999(2)
    2 0.51328(3)
    3 1 0.51858(3) 0.51932
    4 0.52145(5)
    5 0.52319(8)
    6 0.52451(11)
    1 0.51330(3) 0.51932
    2 0.54805(2) 0.55029
    3 2 0.56089(3) 0.56138
    4 0.56758(6) 0.56708
    5 0.57163(6) 0.57056
    6 0.57450(13) 0.57289
    1 0.51856(2) 0.51932
    2 0.56089(5) 0.56138
    3 3 0.57695(5) 0.57682
    4 0.58526(4) 0.58483
    5 0.59047(4) 0.58974
    6 0.59402(10) 0.59305
@@
    5 1 0.66662(5)
    4 1 0.66658(6)
    5 2 0.66679(4)
    4 2 0.66678(2)
3.2.3 Cuboids in 3D :
As an example of a semicontinuum system in 3D, we consider overlapping cuboids in lanes having continuous $X$ and $Y$ coordinates while having discrete $Z$ coordinates (see Fig.3 (a)). The thresholds are determined as in the 2D case and are given in Table. 7 along with the theoretical predictions based on Eq. (8). We can see that, in agreement with predictions of Eq. (8), threshold $\phi_{c}^{k_{1},k_{2},k_{3}}$ is independent of $k_{1}$ and $k_{2}$, for a given value of $k_{3}$. The numerical estimates of the thresholds are obtained by assuming $B_{c}\approx 2.5978$  [20] to be that of the corresponding continuum problem.
From Table 7, we can also infer that the percolation threshold for cubes in the semicontinuum shows an increasing trend with the side-length. Although we have only two data points in Table. 7 corresponding to cubes ($k_{1}=k_{2}=k_{3}=2$ 3), already this is in contrast to the case of cubes on a lattice whose threshold values first show a decrease and then an increase to go to its continuum value of $\approx 0.277$ [20].
$k_{1}$ $k_{2}$ $k_{3}$ $\phi_{c}^{x}$ $\phi_{c}^{y}$ $\phi_{c}^{z}$ $\phi_{c}$ Average $\phi_{c}$ theory (See Eq.(8), $B_{c}\approx 2.5978$ [20])
@@ 4 Conclusion
We analyze a number of percolation problems of overlapping objects in 2D and 3D. In 2D, we considered the percolation of overlapping rectangles, disks, and parallelograms, and in 3D, we considered cuboids. We adapted the excluded volume argument to a semicontinuum setting and studied the dependence of the percolation threshold on the linear measures characterizing the basic percolation units considered. We show that in 2D, for rectangles, the percolation threshold is independent of the length (linear measure along the continuum direction) of the rectangles and is a constant for a given width (linear measure along the discrete direction), even when we have a distribution of lengths. This contrasts with the aspect-ratio-independent thresholds seen in continuum models. Aligned overlapping squares in lanes show a monotonic increase in threshold with increasing size of the squares. This is in contrast to the non-monotonic behavior for the threshold observed for overlapping squares on lattices [20]. We analyzed overlapping disks and parallelograms in 2D semicontinuum geometries. For disks, we find that the threshold decreases monotonically with their radius.
@@
Models in lattice and continuum geometries have been extensively studied in several fields within Science, Engineering, and Mathematics. In many of them, studying how the properties of a system change as we go from a lattice to continuum geometry is a question of great significance. The semicontinuum geometry introduced here describes a system that lies between the two extremes of lattice and continuum geometries. We studied the specific problem of the percolation of overlapping objects in such geometries, but it is obvious that many other models may be considered in such geometries and could be of interest in many fields opening up directions for further investigations. A particular example is that of models with hardcore exclusion that have been extensively studied in lattices and show properties such as jamming and order-disorder transitions [46, 47].
Finally, we note that numerical values of the thresholds obtained here are based on the assumption that the connectivity at criticality is the same for similar shapes in continuum and semicontinuum systems, which allowed us to use the known values of $B_{c}$ from continuum systems. Our theoretical predictions of the numerical values in the models discussed turns out to be very accurate in most cases, justifying the assumption. However, theoretical treatments dedicated to semicontinuum geometries may be required to treat other percolation problems in such settings.
@@
Code availability
Sample source code for our project is available on the GitHub repository.
https://github.com/jasnaCK/Semicontinuum-Percolation